OpenStudy (anonymous):

Can someone help me find the derivative of the following please: y = √3x + 2 cos(x), 0 ≤ x < 2π

OpenStudy (anonymous):

I know I can rewrite it and have \[y=3^{1/2}x+2\cos (x)\]

OpenStudy (anonymous):

\[y'= \frac{ 3 }{ 2 }^{-1/2}x-2\sin x\]

OpenStudy (anonymous):

I think I forgot a step but, I don't know for sure.

OpenStudy (zarkon):

\[\frac{d}{dx}[\sqrt{3}\cdot x]=\sqrt{3}\]

OpenStudy (anonymous):

okay so would I get \[y'=\sqrt{3}-2\sin x\]

OpenStudy (zarkon):

yes

OpenStudy (anonymous):

oh okay I was way over thinking that. Thanks a bunch. I have been working on calculus homework for the past 2 days. I think I need to take a break. I should have known that.